# Label Distribution Learning using the Squared Neural Family on the Probability Simplex

### v1

@@ Abstract
Label distribution learning (LDL) provides a framework wherein a distribution over categories rather than a single category is predicted, with the aim of addressing ambiguity in labeled data. Existing research on LDL mainly focuses on the task of point estimation, i.e., pinpointing an optimal distribution in the probability simplex conditioned on the input sample. In this paper, we estimate a probability distribution of all possible label distributions over the simplex, by unleashing the expressive power of the recently introduced Squared Neural Family (SNEFY). With the modeled distribution, label distribution prediction can be achieved by performing the expectation operation to estimate the mean of the distribution of label distributions. Moreover, more information about the label distribution can be inferred, such as the prediction reliability and uncertainties. We conduct extensive experiments on the label distribution prediction task, showing that our distribution modeling based method can achieve very competitive label distribution prediction performance compared with the state-of-the-art baselines. Additional experiments on active learning and ensemble learning demonstrate that our probabilistic approach can effectively boost the performance in these settings, by accurately estimating the prediction reliability and uncertainties.
@@ 1 Introduction
Label distribution learning (LDL) is a technique which handles ambiguity in multi-class classification, by utilising simplex-valued rather than categorical-valued labels in training data. Unlike traditional multi-class and multi-label learning paradigms, which assign a deterministic label prediction to instances, LDL corresponds to the question “How well does each of the labels describe an instance?”, by using a discrete probability distribution to characterize the instance’s affinity to all candidate categories. For example, when we predict the functionality of a district in a city, we might predict a result such as: the district has 20% functionality for business, 40% functionality for entertainment, and 40% functionality for education.
@@
Many LDL algorithms have been proposed to directly predict label distribution vectors from instance features, by adapting machine learning algorithms designed for “hard” label prediction to the “soft” label prediction setting. Though a discrete distribution among candidate labels is predicted, existing LDL algorithms still operate at the level of point estimation, i.e., they search for a single point on a probability simplex (the set of all possible label distributions) for each given instance. The point estimation paradigm is particularly susceptible to data uncertainty and inexact mappings between instances and labels, due to the inherent complexity of the data collection and generation processes. Therefore, modeling the probability distribution of label distribution vectors, i.e., the probability distribution supported on the probability simplex, is an important step towards trustworthy LDL. An additional bonus of the distribution modeling is the ability to quantify the prediction reliability and uncertainties, which not only facilitates reliable model deployment in real-world safety critical applications, but is also essential to various reliability/uncertainty-aware tasks, like pseudo labeling, active learning, and ensemble learning.
Contributions
@@
In this paper, we propose a novel LDL framework, SNEFY-LDL, by unleashing the probability modeling power of the recently proposed Squared Neural Family (SNEFY) (Tsuchida et al., 2023), a new class of tractable probability models. By restricting the support set of SNEFY to a probability simplex, SNEFY-LDL provides an expressive conditional distribution modeling of the label distribution vector given the input sample. The conditional distribution model has a closed-form normalizing constant, guaranteeing computational tractability. In this way, model parameters can be learned efficiently by maximizing the conditional likelihoods of training samples with stochastic gradient descent. Through computing the expectation of the conditional distribution, we can obtain the closed-form distribution mean, as a way to summarize the fitted model and predict label distributions for unlabeled samples. However, this is only one way to use the fitted model and the probability density values can be used to evaluate the reliability of label distribution predictions.
@@
We conduct extensive experiments on fourteen real-world datasets to evaluate the label distribution prediction performance of the proposed SNEFY-LDL model. Experimental results show that SNEFY-LDL can achieve very competitive label distribution prediction performance, compared with the state-of-the-art LDL baselines. In addition, label distribution active learning and ensemble learning experiments are carried out to verify the efficacy of SNEFY-LDL in quantifying prediction uncertainties. The max-entropy principle is used to achieve active learning with the estimated SNEFY-LDL entropy, i.e., select the most informative unlabeled samples with the largest entropy values, query their labels and augment training samples, to attain the largest performance gain of the re-trained LDL model. Experimental results show that the max-entropy principle achieves significantly better active learning performance than the representativeness based active learning baselines. The experiments on ensemble learning demonstrate that SNEFY-LDL gives a further usecase for the fitted probabilistic model, as it provides an intelligent mechanism for weighting base learners, significantly outperforming the uniform weighting strategy.
@@ 2 Related Work
LDL is first proposed by (Geng et al., 2013) to solve the facial age estimation problem with insufficient training samples. Since then, a series of LDL algorithms have been developed, which can be divided into three categories: Problem Transformation (PT), Algorithm Adaptation (AA) and Specialized Algorithms (SA) (Geng, 2016).
Problem Transformation
@@
PT (Geng, 2016) transforms the LDL problem into the single-label classification problem, by decomposing each training sample assigned with a label distribution into a set of duplicate training samples. Each of them is assigned with a different label and accounts for a partial sample in proportion to the label probability value, and is then used to train the single-label classifiers. The label likelihoods predicted by the single-label classifiers are then aggregated to form the final prediction of label distributions. PT-Bayes (Geng, 2016) and PT-SVM (Geng, 2016; Geng and Hou, 2015) transform the LDL problem into the single-label multi-class classification problem and respectively employ Bayes and SVM as the single-label classifiers. DF-LDL (González et al., 2021a) decomposes the label distribution prediction task into a number of one-versus-one binary classification tasks, and fuses the binary classification likelihoods to form the final label distribution prediction.
Algorithm Adaptation
@@
AA (Geng, 2016) adapts traditional single-label classification models into the LDL setting, by leveraging the models’ compatibility in outputting a soft label distribution vector. Derived from the K Nearest Neighbor (KNN) algorithm (Wu et al., 2008), AA-KNN (Geng, 2016) predicts samples’ label distributions by averaging the label distributions of their $k$ nearest neighbors in feature space. AA-BP (Geng, 2016) constructs a three-layer neural network and adopts the softmax function as the activation of the output layer, making the neural network naturally produce a label distribution for each example. The neural network is trained by minimizing the sum of squared errors between the model output label distributions and the ground-truth label distributions.
Specialized Algorithms
@@
SA (Geng, 2016) designs algorithms from scratch to directly solve the LDL problem. SA-IIS (Geng et al., 2013) and SA-BFGS (Geng, 2016) use the maximum entropy model to parameterize label distributions. They are trained by minimizing the Kullback-Leibler (KL) divergence between the model output and ground-truth label distributions, where Improved Iterative Scaling (IIS) (Malouf, 2002) and BFGS (Nocedal and Wright, 1999) are respectively leveraged by SA-IIS and SA-BFGS as optimizers. CPNN (Geng et al., 2013) uses a neural network to parameterize the joint probability distribution between sample features and labels following Modha’s probability distribution formulation (Modha and Fainman, 1994). BCPNN (Yang et al., 2017) and ACPNN (Yang et al., 2017) then improve on CPNN through leveraging binary label encoding and augmenting training samples respectively. LDLF (Shen et al., 2017) employs differentiable decision trees (Kontschieder et al., 2015) to model label distributions and KL divergence is used to design the learning objective. LDL-SCL (Jia et al., 2021) forces the label distributions of samples located closely in feature space to be similar to each other. LDL-LRR (Jia et al., 2023b) and LDL-DPA (Jia et al., 2023a) maintain the relative importance ranking between different labels in label distribution modeling, by penalizing a label importance ranking loss in their learning objectives.
Extensions
@@
In addition, LDL has been extended to other tasks, like label enhancement (recovering label distributions from the one-hot single-label assignments) (Xu et al., 2019b, 2020; Zheng et al., 2023), multi-class classification (Wang and Geng, 2019, 2021b, 2021a), learning with incomplete supervision (Xu and Zhou, 2017), oversampling (González et al., 2021b), ordinal LDL (Wen et al., 2023), semi-supervised learning (Xie et al., 2023), and label calibration (He et al., 2024). Furthermore, LDL has been applied to solve numerous real-world problems, including facial age estimation (Geng et al., 2013), facial emotion recognition (Chen et al., 2020), head pose estimation (Xu et al., 2019a), crowd opinion prediction (Geng and Hou, 2015), emphasis selection (Shirani et al., 2019), lesion counting (Wu et al., 2019), and urban function distribution (Huang et al., 2023). It is worth noting that there is a related topic in the classical statistics literature, termed compositional data analysis (Greenacre, 2021), with a broad range of applications including geochemistry (e.g. labels correspond to mineral compositions) and ecology (e.g. labels correspond to relative abundance of species).
@@ 3 Problem Definition
Assume we are given a set of $N$ i.i.d training samples $\mathcal{X}=\{\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{N}\}$ with each sample $\bm{x}\in\mathcal{X}$ located in the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. In addition, each sample $\bm{x}\in\mathcal{X}$ is described by a $L$-dimensional label distribution vector $\bm{\ell}_{\bm{x}}\in\mathbb{R}^{L}$ that takes values in the $(L-1)$-simplex $\Delta^{L-1}$, corresponding to a set of $L$ given labels $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{L}\}$. The $l$-th entry $\bm{\ell}_{\bm{x}}^{y_{l}}$ of the label distribution vector $\bm{\ell}_{\bm{x}}$ corresponds to the probability of sample $\bm{x}$ belonging to the $l$-th class $y_{l}$, i.e., $\bm{\ell}_{\bm{x}}^{y_{l}}=\mathrm{P}(y_{l}|\bm{x})\geq 0$, satisfying the constraint that $\sum_{l=1}^{L}\bm{\ell}_{\bm{x}}^{y_{l}}=1$.
@@
For any sample $\bm{x}\in\mathbb{R}^{d}$ with an unknown label distribution, one can associate to it a label distribution vector $\bm{\ell}^{*}_{\bm{x}}$ obtained as the mean of the fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x})$:
$\bm{\ell}^{*}_{\bm{x}}=\int_{\Delta^{L-1}}\bm{\ell}\,\mathrm{P}(d\bm{\ell}|\bm{x}),$ (1)
but the fitted model $\mathrm{P}(d\bm{\ell}|\bm{x})$ can serve many other purposes.
@@ 4 Preliminaries on SNEFY
where $\bm{t}(\cdot):\mathcal{Z}\rightarrow\mathbb{R}^{D}$ is the sufficient statistic, $\sigma$ is the activation function, $\bm{W}\in\mathbb{R}^{n\times D}$ and $\bm{b}\in\mathbb{R}^{n}$ are respectively the weight matrix and bias vector at the hidden layer of neural network $\bm{f}$, and $\bm{V}\in\mathbb{R}^{m\times n}$ are $\bm{f}$’s readout parameters. $\bm{\Theta}=(\bm{W},\bm{b})\in\mathbb{R}^{n\times(D+1)}$ denotes the concatenation of $\bm{W}$ and $\bm{b}$ and its $i$th row is denoted as $\bm{\theta}_{i}=(\bm{w}_{i},b_{i})\in\mathbb{R}^{D+1}$.
@@
Under varying choices of the activation function $\sigma$, sufficient statistic $\bm{t}$, and the base measure $\mu$, $\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})$ is able to be computed in closed form (see Table 1 of (Tsuchida et al., 2023)) in $\mathcal{O}(D)$. This makes SNEFY a tractable probability distribution model, with great expressivity and computational efficiency.
@@
(Tsuchida et al., 2023) Let $\mathtt{z}=(\mathtt{z}_{1},\mathtt{z}_{2})$ jointly follow a SNEFY distribution, with support set $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}$, sufficient statistic $\bm{t}$, activation function $\sigma$, base measure $\mu$, as well as parameters $\bm{V}$ and $\bm{\Theta}=([\bm{W}_{1},\bm{W}_{2}],\bm{b})$. Assume that $\mu(d\bm{z})=\mu_{1}(d\bm{z}_{1})\mu_{2}(d\bm{z}_{2})$ and $\bm{t}(\bm{z})=(\bm{t}_{1}(\bm{z}_{1}),\bm{t}_{2}(\bm{z}_{2}))$. Then the conditional distribution of $\mathtt{z}_{1}$ given $\mathtt{z}_{2}=\bm{z}_{2}$ is also a SNEFY distribution with support set $\mathcal{Z}_{1}$, sufficient statistic $\bm{t}_{1}$, activation function $\sigma$, base measure $\mu_{1}$, as well as parameters $\bm{V}$ and $\bm{\Theta}_{1|2}=(\bm{W}_{1},\bm{W}_{2}\bm{t}_{2}(\bm{z}_{2})+\bm{b})$.

### v2

@@ Abstract
Label distribution learning (LDL) provides a framework wherein a distribution over categories rather than a single category is predicted, with the aim of addressing ambiguity in labeled data. Existing research on LDL mainly focuses on the task of point estimation, i.e., finding an optimal distribution in the probability simplex conditioned on the given sample. In this paper, we propose a novel label distribution learning model SNEFY-LDL, which estimates a probability distribution of all possible label distributions over the simplex, by unleashing the expressive power of the recently introduced Squared Neural Family (SNEFY), a new class of tractable probability models. As a way to summarize the fitted model, we derive the closed-form label distribution mean, variance and covariance conditioned on the given sample, which can be used to predict the ground-truth label distributions, construct label distribution confidence intervals, and measure the correlations between different labels. Moreover, more information about the label distribution prediction uncertainties can be acquired from the modeled probability density function. Extensive experiments on conformal prediction, active learning and ensemble learning are conducted, verifying SNEFY-LDL’s great effectiveness in LDL uncertainty quantification. The source code of this paper is available at https://github.com/daokunzhang/SNEFY-LDL.
@@ 1 Introduction
Label distribution learning (LDL) is a technique which handles ambiguity in multi-class classification, by utilizing simplex-valued rather than categorical-valued labels in training data. Unlike traditional multi-class and multi-label learning paradigms, which assign a deterministic label prediction to instances, LDL corresponds to the question "How well does each of the labels describe an instance?", by using a discrete probability distribution to characterize each label’s composition ratio in jointly describing the given instance. For example, when we predict the functionality of a district in a city, we might predict a result such as: the district has 20% functionality for business, 40% functionality for entertainment, and 40% functionality for education.
@@
Many LDL algorithms have been proposed to directly predict label distribution vectors from instance features, by adapting machine learning algorithms designed for "hard" label prediction to the "soft" label prediction setting. Though a discrete distribution among candidate labels is predicted, existing LDL algorithms still operate at the level of point estimation, i.e., they search for a single point on a probability simplex (the set of all possible label distributions) for each given instance. The point estimation paradigm is particularly susceptible to data uncertainty and inexact mappings between instances and labels, due to the inherent complexity of the data collection and generation processes. Therefore, modeling the probability distribution of label distribution vectors, i.e., the probability distribution supported on the probability simplex, is an important step towards trustworthy LDL. An additional bonus of the distribution modeling is the ability to quantify the prediction reliability and uncertainties, which not only facilitates reliable model deployment in real-world safety critical applications, but is also essential to various reliability/uncertainty-aware tasks, like pseudo labeling, active learning, and ensemble learning.
@@
Contributions. In this paper, we propose a novel LDL framework, SNEFY-LDL, by unleashing the probability modeling power of the recently introduced Squared Neural Family (SNEFY) [Tsuchida et al., 2023], a new class of tractable probability models. By restricting the support set of SNEFY to a probability simplex, SNEFY-LDL constructs an expressive multimodal distribution modeling of the label distribution vector conditioned the given sample. The conditional distribution model has a closed-form normalizing constant, guaranteeing computational tractability. In this way, model parameters can be learned efficiently by maximizing the conditional likelihoods of training samples with stochastic gradient descent. As a way to summarize the fitted model, we derive the closed-form label distribution mean, variance and covariance conditioned on the given sample, which can be used to predict the ground-truth label distribution, construct label distribution confidence intervals, and measure the correlations between different labels. However, the fitted model is not limited to these usecases, and the probability density values can be used to directly evaluate the reliability of label distribution predictions.
@@
We conduct extensive experiments on label distribution conformal prediction, active learning and ensemble learning to verify the efficacy of SNEFY-LDL in quantifying prediction uncertainties. For the task of conformal prediction, we use the SNEFY-LDL’s closed-form conditional mean and variance of label distribution predictions to construct a confidence interval for each label’s composition ratio in describing the given instance and calibrate the confidence interval through conformal prediction [Angelopoulos and Bates, 2021]. Experimental results show that the confidence intervals constructed by the SNEFY-LDL model have greater adaptivity than the confidence intervals constructed by the naive Dirichlet distribution. The max-entropy principle is used to achieve active learning with the estimated SNEFY-LDL entropy, i.e., select the most informative unlabeled samples with the largest entropy values, query their labels and augment training samples, to attain the largest performance gain of the re-trained LDL model. Experimental results show that the max-entropy principle achieves significantly better active learning performance than the representativeness based active learning baselines. The experiments on ensemble learning demonstrate that SNEFY-LDL gives a further usecase for the fitted probabilistic model, as it provides an intelligent mechanism for weighting base learners, significantly outperforming the uniform weighting strategy.
@@ 2 Related Work
LDL is first proposed by Geng et al. [2013] to solve the facial age estimation problem. Since then, a series of LDL algorithms have been developed, which are mainly in three categories: Problem Transformation (PT), Algorithm Adaptation (AA) and Specialized Algorithms (SA) [Geng, 2016].
@@
Problem Transformation. PT [Geng, 2016] transforms the LDL problem into the single-label classification problem, by decomposing each training sample assigned with a label distribution into a set of duplicate training samples. Each of them is assigned with a different label and accounts for a partial sample in proportion to the label probability value, and is then used to train the single-label classifiers. The label likelihoods predicted by the single-label classifiers are then aggregated to form the final prediction of label distributions. PT-Bayes [Geng, 2016] and PT-SVM [Geng and Hou, 2015] transform the LDL problem into the single-label multi-class classification problem and respectively employ Bayes and SVM as the single-label classifiers. DF-LDL [González et al., 2021a] decomposes the label distribution prediction task into a number of one-versus-one binary classification tasks, and fuses the binary classification likelihoods to form the final label distribution predictions.
@@
Algorithm Adaptation. AA [Geng, 2016] adapts traditional single-label classification models into the LDL setting, by leveraging the models’ compatibility in outputting a soft label distribution vector. Derived from the K Nearest Neighbor (KNN) algorithm [Wu et al., 2008], AA-KNN [Geng, 2016] predicts samples’ label distributions by averaging the label distributions of their $k$ nearest neighbors in feature space. AA-BP [Geng, 2016] constructs a three-layer neural network and adopts the softmax function as the activation of the output layer, making the neural network naturally produce a label distribution for each example. The neural network is trained by minimizing the sum of squared errors between the model-output label distributions and the ground-truth label distributions.
@@
Specialized Algorithms. SA [Geng, 2016] designs algorithms from scratch to directly solve the LDL problem. SA-IIS [Geng et al., 2013] and SA-BFGS [Geng, 2016] use the maximum entropy model to parameterize label distributions. They are trained by minimizing the Kullback-Leibler (KL) divergence between the model-output and ground-truth label distributions, where Improved Iterative Scaling (IIS) [Malouf, 2002] and BFGS [Nocedal and Wright, 1999] are respectively leveraged by SA-IIS and SA-BFGS as optimizers. CPNN [Geng et al., 2013] uses a neural network to parameterize the joint probability distribution between sample features and labels following Modha’s probability distribution formulation [Modha and Fainman, 1994]. BCPNN [Yang et al., 2017] and ACPNN [Yang et al., 2017] then improve on CPNN through leveraging binary label encoding and augmenting training samples respectively. LDLF [Shen et al., 2017] employs differentiable decision trees [Kontschieder et al., 2015] to model label distributions and KL divergence is used to design the learning objective. LDL-SCL [Jia et al., 2021] forces the label distributions of samples located closely in feature space to be similar to each other. LDL-LRR [Jia et al., 2023b] and LDL-DPA [Jia et al., 2023a] maintain the relative importance ranking between different labels in label distribution modeling, by penalizing a label importance ranking loss in their learning objectives.
@@
Extensions. In addition, LDL has been extended to other tasks, like label enhancement  [Xu et al., 2019b, 2020, Zheng et al., 2023], multi-class classification [Wang and Geng, 2019, 2021b, 2021a], learning with incomplete supervision [Xu and Zhou, 2017], oversampling [González et al., 2021b], ordinal LDL [Wen et al., 2023], semi-supervised learning [Xie et al., 2023], and label calibration [He et al., 2024]. Furthermore, LDL has been applied to solve numerous real-world problems, including facial age estimation [Geng et al., 2013], facial emotion recognition [Chen et al., 2020], head pose estimation [Xu et al., 2019a], crowd opinion prediction [Geng and Hou, 2015], emphasis selection [Shirani et al., 2019], lesion counting [Wu et al., 2019], and urban functionality prediction [Huang et al., 2023]. It is worth noting that there is a related topic in the classical statistics literature, termed compositional data analysis [Greenacre, 2021], with a broad range of applications including geochemistry (e.g. labels correspond to mineral compositions) and ecology (e.g. labels correspond to relative abundance of species).
@@ 3 Problem Definition
Assume we are given a set of $N$ i.i.d training samples $\mathcal{X}=\{\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{N}\}$ with each sample $\bm{x}\in\mathcal{X}$ located in the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. In addition, each sample $\bm{x}\in\mathcal{X}$ is described by a $L$-dimensional label distribution vector $\bm{\ell}_{\bm{x}}\in\mathbb{R}^{L}$ that takes values in the $(L-1)$-simplex $\Delta^{L-1}$, corresponding to a set of $L$ given labels $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{L}\}$. The $l$-th entry $\bm{\ell}_{\bm{x}}^{y_{l}}$ of the label distribution vector $\bm{\ell}_{\bm{x}}$ corresponds to the composition ratio of the $l$-th label $y_{l}$ in describing $\bm{x}$, satisfying the constraint that $\sum_{l=1}^{L}\bm{\ell}_{\bm{x}}^{y_{l}}=1$.
@@ 4 Preliminaries on SNEFY
where $\bm{t}(\cdot):\mathcal{Z}\rightarrow\mathbb{R}^{D}$ is the sufficient statistic, $\sigma$ is the activation function, $\bm{W}\in\mathbb{R}^{n\times D}$ and $\bm{b}\in\mathbb{R}^{n}$ are respectively the weight matrix and bias vector at the hidden layer of neural network $\bm{f}$, and $\bm{V}\in\mathbb{R}^{m\times n}$ are $\bm{f}$’s readout parameters. $\bm{\Theta}=(\bm{W},\bm{b})\in\mathbb{R}^{n\times(D+1)}$ is the concatenation of $\bm{W}$ and $\bm{b}$ and its $i$th row is denoted as $\bm{\theta}_{i}=(\bm{w}_{i},b_{i})\in\mathbb{R}^{D+1}$, where $\bm{w}_{i}\in\mathbb{R}^{D}$ is the $i$th row of $\bm{W}$ and $b_{i}$ is the $i$th element of $\bm{b}$.
@@
Under varying choices of the activation function $\sigma$, sufficient statistic $\bm{t}$, and the base measure $\mu$, $\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})$ is able to be computed in closed form (see Table 1 of [Tsuchida et al., 2023]) in $\mathcal{O}(D)$. This makes SNEFY a tractable probability distribution model, with great expressivity and computational efficiency.
@@
[Tsuchida et al., 2023] Let $\mathtt{z}=(\mathtt{z}_{1},\mathtt{z}_{2})$ jointly follow a SNEFY distribution, with support set $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}$, sufficient statistic $\bm{t}$, activation function $\sigma$, base measure $\mu$, as well as parameters $\bm{V}$ and $\bm{\Theta}=([\bm{W}_{1},\bm{W}_{2}],\bm{b})$. Assume that $\mu(d\bm{z})=\mu_{1}(d\bm{z}_{1})\mu_{2}(d\bm{z}_{2})$ and $\bm{t}(\bm{z})=(\bm{t}_{1}(\bm{z}_{1}),\bm{t}_{2}(\bm{z}_{2}))$. Then the conditional distribution of $\mathtt{z}_{1}$ given $\mathtt{z}_{2}=\bm{z}_{2}$ is also a SNEFY distribution with support set $\mathcal{Z}_{1}$, sufficient statistic $\bm{t}_{1}$, activation function $\sigma$, base measure $\mu_{1}$, as well as parameters $\bm{V}$ and $\bm{\Theta}_{1|2}=(\bm{W}_{1},\bm{W}_{2}\bm{t}_{2}(\bm{z}_{2})+\bm{b})$.